OpenStudy (anonymous):

what is 2-11x^2-8x+6x^2

OpenStudy (catlover5925):

\(\huge\color{lime}{Welcome~To~Openstudy!}\)

OpenStudy (anonymous):

i think it is 2

OpenStudy (anonymous):

what is the first step here?

OpenStudy (anonymous):

x = {2.3763293357, -1.5429960023}

OpenStudy (anonymous):

Let's simplify step-by-step. 2−11x2−8x+6x2 =2+−11x2+−8x+6x2

OpenStudy (anonymous):

Combine Like Terms:

OpenStudy (anonymous):

=2+−11x2+−8x+6x2 =(−11x2+6x2)+(−8x)+(2) =−5x2+−8x+2

OpenStudy (anonymous):

i think cookinggeek is right

OpenStudy (anonymous):

Answer: =−5x2−8x+2